OpenStudy (anonymous):

Solve using the addition method -3x+y=-2 -28x+4y=8

OpenStudy (britanyashton):

are you solving for x and y? addition method of what?

OpenStudy (anonymous):

yes im solving for x and y and i meant elimination method

OpenStudy (anonymous):

@britanyashton

OpenStudy (anonymous):

@BunnyBree can you help?

OpenStudy (britanyashton):

oh. okay. i get this. If bunnyBee doesnt know, then ill help.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Okay, so if this if for the elimination method, we need to eliminate one of the variables to solve for another. So, eliminating y would be easier than eliminating x. Multiply the first equation by -4. (-3x+y=-2) x -4 -28x+4y=8 We get: 12x -4y = 8 -28x + 4y = 8 See how -4y and 4y cancel each other? Now we have: 12x = 8 -28x = 8 Add. -16x = 16 Divide to isolate x. x = -1 <---- One of the solutions. Now just plug x into one of the original equations to obtain y.

OpenStudy (anonymous):

would y = -6?

OpenStudy (anonymous):

I'm going to just choose the first equation and plug x in. -3x+y=-2 -3(-1) + y = -2 3 + y = -2 Subtract 3 from both sides. y = -5 So, your solutions are: x = -1, and y = -5. (:

OpenStudy (anonymous):

Do you understand how I got -5 for y?

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

You're welcome. (: